# Dust-obscured Galaxies with Broken Power-law Spectral Energy Distributions Discovered by UNIONS

### v1

@@ 1 Introduction
To study dusty populations of galaxies and AGNs, we focus on dust-obscured galaxies (DOGs: Dey et al. 2008). They are selected based on their extremely red color in optical to mid-infrared (MIR) wavelengths in the observed frame, which is caused by extreme dust extinction in optical and re-emission in infrared (IR). DOGs have been divided into two subclasses based on their optical-MIR spectral energy distributions (SEDs), which are known as “bump DOGs” and “power-law (PL) DOGs” (e.g., Dey et al. 2008; Melbourne et al. 2012; Toba et al. 2015; Suleiman et al. 2022; Yu et al. 2024). Bump DOGs are characterized by a bump feature (corresponding to rest-frame 1.6 $\mu$m bump) around near-infrared (NIR) in their SEDs mainly caused by the stellar photospheres of cool stars, which suggests that star-forming activity dominates those DOGs (e.g., Desai et al. 2009; Melbourne et al. 2012; Yu et al. 2024; see also Pope et al. 2008; Riguccini et al. 2011). PL DOGs are characterized by power-law shapes of their SEDs from optical to MIR wavelengths, and represent dusty AGNs (e.g., Bussmann et al. 2009; Melbourne et al. 2012; Zou et al. 2020; Yu et al. 2024; see also Alonso-Herrero et al. 2006; Imanishi et al. 2010; Ichikawa et al. 2014). The presence of obscured AGNs in PL DOGs is also supported by X-ray studies, showing that the majority of PL DOGs have a large column density ($N_{\mathrm{H}}>10^{22}$ cm-2; e.g., Lanzuisi et al. 2009; Corral et al. 2016; Riguccini et al. 2019; Toba et al. 2020; Zou et al. 2020; Kayal & Singh 2024, Noboriguchi et al. in preparation).
@@
In summary, researching DOGs and their subclasses is effective in understanding the evolutionary stage of major mergers of gas-rich galaxies. Constructing a large sample of DOGs is challenging because they are spatially rare and significantly faint in optical wavelengths, requiring a wide survey area and deep sensitivity. Previous studies have identified up to several thousand DOGs using wide-area surveys or deep fields (e.g., Toba & Nagao 2016; Yu et al. 2024). However, most of DOGs found in those earlier surveys have not been classified into subclasses (bump and PL DOGs) due to the lack of deep multi-band photometry, particularly in the NIR. So far, only a few hundred DOGs have been classified in previous studies, which is not sufficient to study the statistical properties for each subclass of DOGs (e.g., luminosity function). To address this, we focus on the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS; Gwyn et al. 2025) and $Euclid$ ($Euclid$ Collaboration 2024). UNIONS is an ongoing multi-band optical imaging survey aiming to obtain deep $ugriz$ images over $\sim$ 5,000 deg2 of the northern sky, which mostly overlaps with the northern part of the $Euclid$ survey area. $Euclid$ is a space telescope that is currently carrying out single-band optical and multi-band NIR imaging survey with high sensitivity. By combining UNIONS and $Euclid$ data, we will be able to construct a statistically large DOG sample with deep multi-band photometry in the near future. This will allow us to study statistical properties for each subclass in detail.
@@ 2.4 Selection of DOGs
where $i$ and [22] represent the magnitudes in the UNIONS $i$-band and WISE 22 $\mu$m band, respectively (Toba et al. 2015; see also Dey et al. 2008). Using this criterion, we successfully identified 382 DOGs over $\sim$ 170 deg2. However, we excluded 6 objects that were undetected at both 4.6 and 12 $\mu$m, as these could not be classified into any subclasses (see Section 3.2). Therefore, we focus on the remaining 376 DOGs in the following analyses. Note that we use only the $u$, $g$, $r$, and $i$-band UNIONS photometry to study their SEDs, because the $z$-band photometry of UNIONS2000 is not currently available in the overlap region of UNIONS and UKIDSS.
@@ 3.3 Comparison of slopes
Figure 3 shows the difference between the optical-NIR and MIR slopes ($\alpha_{\rm optNIR}-\alpha_{\rm MIR}$) of PL DOGs. The mean and standard deviation of $\alpha_{\rm optNIR}-\alpha_{\rm MIR}$ for 244 PL DOGs are 1.0 and 0.7, respectively. The distribution of $\alpha_{\rm optNIR}-\alpha_{\rm MIR}$ appears to have a bimodal structure, suggesting the possible presence of two populations corresponding to NPL and BPL DOGs. Based on this, we define BPL DOGs as PL DOGs with $\alpha_{\rm optNIR}-\alpha_{\rm MIR}\geq 1.0$, which corresponds to the local minimum between the two potential peaks. As a result, roughly half (120/244) of the PL DOGs are classified as BPL DOGs. This large fraction of BPL DOGs among PL DOGs possibly reflects a selection bias, which will be discussed in Section 4.2. We show the individual and median SEDs of NPL and BPL DOGs in Figure 4.
@@ 4.3 Relationship with other dusty AGNs
Wu et al. (2012) and Tsai et al. (2015) compared the SEDs of Hot DOGs with those of torus model (Polletta et al. 2006, 2007; see Section 4.1.2) and found that the slopes of Hot DOGs are steeper than those of the torus SED model in the rest-frame optical wavelength, suggesting that Hot DOGs are more heavily obscured than the torus model. In Section 4.1.1, we showed that BPL DOGs have a similar $\alpha_{\rm optNIR}$ to the torus SED model, suggesting that Hot DOGs are the most heavily obscured AGNs among various DOG populations.
Figure 8: Comparison of the SEDs of broken-PL DOGs and the stacked SEDs of LRDs (Pérez-González et al. 2024; Akins et al. 2024; see also Li et al. 2025) shifted to $z=1$, and normalized at $\lambda_{\rm obs}\sim 0.8$ $\mu$m. The texts and number in parentheses indicate the reference papers and the number of objects in our BPL DOG sample.
@@ acknowledgments
This work was financially supported by JSPS KAKENHI grants: Nos. 20H01949 (T.N.), 23H01215 (T.N.), 23K22537 (Y.T.) and 24K00684 (M.O.). H. H. is supported by a DFG Heisenberg grant (Hi 1495/5-1), the DFG Collaborative Research Center SFB1491, an ERC Consolidator Grant (No. 770935), and the DLR project 50QE2305. Y.Z. is supported by JSPS Research Fellowship for Young Scientists.

### v2

@@ 1 Introduction
To study dusty populations of galaxies and AGNs, we focus on dust-obscured galaxies (DOGs: Dey et al. 2008). They are selected based on their extremely red color in optical to mid-infrared (MIR) wavelengths in the observed frame, which is caused by extreme dust extinction in optical and re-emission in infrared (IR). DOGs have been divided into two subclasses based on their optical-MIR spectral energy distributions (SEDs), which are known as “bump DOGs” and “power-law (PL) DOGs” (e.g., Dey et al. 2008; Melbourne et al. 2012; Toba et al. 2015; Suleiman et al. 2022; Yu et al. 2024). Bump DOGs are characterized by a bump feature (corresponding to rest-frame 1.6 $\mu$m bump) around near-infrared (NIR) in their SEDs mainly caused by the stellar photospheres of cool stars, which suggests that star-forming activity dominates those DOGs (e.g., Desai et al. 2009; Melbourne et al. 2012; Yu et al. 2024; see also Pope et al. 2008; Riguccini et al. 2011). PL DOGs are characterized by power-law shapes of their SEDs from optical to MIR wavelengths, and represent dusty AGNs (e.g., Bussmann et al. 2009; Melbourne et al. 2012; Zou et al. 2020; Yu et al. 2024; see also Alonso-Herrero et al. 2006; Imanishi et al. 2010; Ichikawa et al. 2014). The presence of obscured AGNs in PL DOGs is also supported by X-ray studies, showing that the majority of PL DOGs have a large column density ($N_{\mathrm{H}}>10^{22}$ cm-2; e.g., Lanzuisi et al. 2009; Corral et al. 2016; Riguccini et al. 2019; Toba et al. 2020; Zou et al. 2020; Kayal & Singh 2024; Noboriguchi et al. 2025).
@@
In summary, researching DOGs and their subclasses is effective in understanding the evolutionary stage of major mergers of gas-rich galaxies. Constructing a large sample of DOGs is challenging because they are spatially rare and significantly faint in optical wavelengths, requiring a wide survey area and deep sensitivity. Previous studies have identified up to several thousand DOGs using wide-area surveys or deep fields (e.g., Toba & Nagao 2016; Yu et al. 2024, Noboriguchi et al. 2025). However, most of DOGs found in those earlier surveys have not been classified into subclasses (bump and PL DOGs) due to the lack of deep multi-band photometry, particularly in the NIR. So far, only a few hundred DOGs have been classified in previous studies, which is not sufficient to study the statistical properties for each subclass of DOGs (e.g., luminosity function). To address this, we focus on the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS; Gwyn et al. 2025) and $Euclid$ ($Euclid$ Collaboration 2024). UNIONS is an ongoing multi-band optical imaging survey aiming to obtain deep $ugriz$ images over $\sim$ 5,000 deg2 of the northern sky, which mostly overlaps with the northern part of the $Euclid$ survey area. $Euclid$ is a space telescope that is currently carrying out single-band optical and multi-band NIR imaging survey with high sensitivity. By combining UNIONS and $Euclid$ data, we will be able to construct a statistically large DOG sample with deep multi-band photometry in the near future. This will allow us to study statistical properties for each subclass in detail.
@@ 2.4 Selection of DOGs
where $i$ and [22] represent the magnitudes in the UNIONS $i$-band and WISE 22 $\mu$m band, respectively (Toba et al. 2015; see also Dey et al. 2008) 333 Note that the selection criterion defined by Toba et al. (2015) does not take a redshift dependence into account. Actually, some authors reported local DOGs by considering rest-frame DOG color (e.g., Hwang & Geller 2013). On the other hand, our criterion is optimized for DOG search with HSC and WISE, where traditional DOGs in a wide redshift range down to $z\sim 1$ also satisfy the criterion (e.g., Toba et al. 2015; Noboriguchi et al. 2019).. Using this criterion, we successfully identified 382 DOGs over $\sim$ 170 deg2. However, we excluded 6 objects that were undetected at both 4.6 and 12 $\mu$m, as these could not be classified into any subclasses (see Section 3.2). Therefore, we focus on the remaining 376 DOGs in the following analyses. Note that we used only the $u$, $g$, $r$, and $i$-band UNIONS photometry to study their SEDs, because the $z$-band photometry of UNIONS2000 is not currently available in the overlap region of UNIONS and UKIDSS.
@@ 3.3 Comparison of slopes
Figure 3 shows the difference between the optical-NIR and MIR slopes ($\alpha_{\rm optNIR}-\alpha_{\rm MIR}$) of PL DOGs. The mean and standard deviation of $\alpha_{\rm optNIR}-\alpha_{\rm MIR}$ for 244 PL DOGs are 1.0 and 0.7, respectively. The distribution of $\alpha_{\rm optNIR}-\alpha_{\rm MIR}$ appears to have a bimodal structure, suggesting the possible presence of two populations corresponding to NPL and BPL DOGs. Based on this, we define BPL DOGs as PL DOGs with $\alpha_{\rm optNIR}-\alpha_{\rm MIR}\geq 1.0$, which corresponds to the local minimum between the two potential peaks. As a result, roughly half (120/244) of the PL DOGs are classified as BPL DOGs. This large fraction of BPL DOGs among PL DOGs possibly reflects a selection bias, which will be discussed in Section 4.2. We show the individual and median SEDs of NPL and BPL DOGs in Figure 4. This figure shows that $\alpha_{\rm MIR}$ is not significantly different between NPL and BPL DOGs, while BPL DOGs show steeper $\alpha_{\rm optNIR}$ than NPL DOGs.
@@
Finally, we performed Kolmogorov–Smirnov test for $\alpha_{\rm optNIR}$ and $\alpha_{\rm MIR}$ between each pair of subclasses (Table 1). The $p$-values are significantly small for all cases, indicating that the distributions of the subclasses in the $\alpha_{\rm optNIR}$-$\alpha_{\rm MIR}$ plane are statistically different with $>3\sigma$ significance.
Table 1: The $p$-value of Kolmogorov–Smirnov test for $\alpha_{\rm optNIR}$ and $\alpha_{\rm MIR}$ between each pair of subclasses.
Pair of subclasses $\alpha_{\rm optNIR}$ $\alpha_{\rm MIR}$
bump vs normal-PL $9.1\times 10^{-13}$ $1.4\times 10^{-30}$
bump vs broken-PL $2.4\times 10^{-4}$ $8.0\times 10^{-64}$
normal-PL vs broken-PL $9.3\times 10^{-26}$ $3.5\times 10^{-16}$
@@ 4.3 Relationship with other dusty AGNs
Wu et al. (2012) and Tsai et al. (2015) compared the SEDs of Hot DOGs with those of torus model (Polletta et al. 2006, 2007; see Section 4.1.2) and found that the slopes of Hot DOGs are steeper than those of the torus SED model in the rest-frame optical wavelength, suggesting that Hot DOGs are more heavily obscured than the torus model. In Section 4.1.1, we showed that BPL DOGs have a similar $\alpha_{\rm optNIR}$ to the torus SED model, suggesting that Hot DOGs are the most heavily obscured AGNs among various DOG populations. In Figure 8, we compare the SEDs of our BPL DOGs with the median SED of Hot DOGs presented by Tsai et al. (2015). For a fair comparison, we shifted the median SED of Hot DOGs from their typical redshift of $z=3.0$ (Tsai et al. 2015) to $z=1.0$, which is more representative of IR-bright DOGs. Figure 8 shows a steeper slope of Hot DOG SED around rest-frame 1–5 $\mu$m range than those of BPL DOG SED, suggesting their heavier obscuration. The flatter SEDs of Hot DOGs in both shorter and longer wavelengths than those of BPL DOGs are due to their stellar emission and the excess of hot dust emission, respectively (e.g., Assef et al. 2015; Tsai et al. 2015).
Figure 8: Comparison of the SEDs of broken-PL DOGs, the median SED of Hot DOGs (Tsai et al. 2015), and the stacked SEDs of LRDs (Pérez-González et al. 2024; Akins et al. 2024; see also Li et al. 2025). The SEDs of Hot DOGs and LRDs are shifted to $z=1$, and normalized at $\lambda_{\rm obs}\sim 0.8$ $\mu$m. The texts and number in parentheses indicate the reference papers and the number of objects in our broken-PL DOG sample.
@@ acknowledgments
The authors gratefully acknowledge the anonymous referee for a careful reading of the manuscript and very helpful comments.
@@
This work was financially supported by JSPS KAKENHI grants: Nos. 20H01949 (T.N.), 23H01215 (T.N.), 23K22537 (Y.T.) and 24K00684 (M.O.). H. H. is supported by a DFG Heisenberg grant (Hi 1495/5-1), the DFG Collaborative Research Center SFB1491, an ERC Consolidator Grant (No. 770935), and the DLR project 50QE2305. Y.Z. is supported by JSPS Research Fellowship for Young Scientists. This work was supported by JSPS Core-to-Core Program (grant number: JPJSCCA20210003).